An airplane went 960 miles in the same length of time that a second plane went 1,200 miles. The speed of the second plane was 40 miles an hour faster than that of the first. How long did each plane travel?
Answer by lwsshak3(11628)

An airplane went 960 miles in the same length of time that a second plane went 1,200 miles. The speed of the second plane was 40 miles an hour faster than that of the first. How long did each plane travel?
***
let x=speed of 1st plane
x+40=speed of 2nd plane
distance/speed=travel time
..
960/x=1200/(x+40)
1200x=960x+38400
240x=38400
x=160
x+40=200
960/160=6
How long did each plane travel? 6 hrs